OpenStudy (anonymous):

If the dimensions of a cylinder are doubled, then its volume is quadrupled. True or False. Please HELP!!!

OpenStudy (anonymous):

8 times

OpenStudy (anonymous):

false

OpenStudy (anonymous):

Hmm... Is there anyway I can check?

OpenStudy (anonymous):

That is the surface area, not counting top and bottom.

OpenStudy (anonymous):

okay, yeah that makes since...

OpenStudy (anonymous):

oh sorry for that blunder

OpenStudy (anonymous):

yeah i just checked it is false. :)

OpenStudy (anonymous):

\[V=\pi r ^{2}h\] for a right circular cylinder

OpenStudy (anonymous):

Yup. Thanks for the help! :)

OpenStudy (anonymous):

If you double the radius, volume is quadrupled. If you double the height, volume is doubled. If you double both... 8 times.

OpenStudy (anonymous):

old volume= (pi)r^2h new volume=(pi)(2r)^2(2h)=8times old volume

OpenStudy (anonymous):

Perfect! Octupled :)